OpenStudy (anonymous):

PLEASE HELP? 1.Write the slope-intercept form of the equation of the line parallel to y = 5x + 2, which passes through the point (4, 9). 2.Write the equation y - 2 = 4(x + 5) in standard form.

OpenStudy (anonymous):

y = 5x + 2....the slope in this equation is 5. A parallel line will have the same slope. now we will use y = mx + b and solve for b slope(m) = 5 (4,9) x = 4 and y = 9 now we sub y = mx + b 9 = 5(4) + b 9 = 20 + b 9 - 20 = b -11 = b your parallel equation is : y = 5x - 11 ========================== y - 2 = 4(x + 5) y - 2 = 4x + 20 y = 4x + 20 + 2 y = 4x + 22 (subtract 4x) -4x + y = 22 (multiply by -1 to make x positive) 4x - y = -22 <-- standard form